# A Practical Review of Mechanistic Interpretability for Transformer-Based Language Models

### v2

@@ 1 Introduction
Related Survey Format Technical Focus Findings Applications Discussion of Future Research
Zhao et al. (2024) Technique-centric Non MI-focused Explanation methods comparison Not a focus Explainability-focused
Ferrando et al. (2024) Technique-centric MI-focused LM properties Not a focus Brief
Bereska & Gavves (2024) Technique-centric MI-focused LM properties AI safety-focused Comprehensive
Ours Task-centric MI-focused LM properties Broad LM applications Comprehensive
Table 1: Comparison of our survey with others on LM interpretability. Our survey brings a unique perspective of task-centric survey and includes a comprehensive review of MI techniques, findings, and applications.
@@
Observing these promises and challenges, we aim to provide a comprehensive review of MI in its applications to interpret transformer-based LMs. In particular, our survey offers a “task-centric” perspective, organizing the taxonomy of MI research with specific research questions or tasks and enabling readers new to the field to quickly identify problems of their interest. Specifically, this is achieved by structuring our taxonomy as a Beginner’s Roadmap to MI (Section 5), where each category and subcategory are centered on overarching research questions and accompanied by actionable workflows. These workflows outline key steps and applicable techniques, guiding practitioners through the entire process from problem formulation to practical implementation. In addition, we discuss the pros and cons of applicable techniques (e.g., vocabulary project methods [nostalgebraist 2020] vs. sparse autonecoders [Bricken et al. 2023]) and present readers with case studies of how prior works fit within the workflow, aiming to aid them in understanding the practices and choosing the best-fit approaches to meet their need. By doing so, our survey serves as a friendly and practical guide for beginning researchers (e.g., new doctoral students) to quickly grasp the field of MI, and for experienced practitioners to more easily organize their study plan.
There exist surveys on relevant topics, but they differ from ours in a number of key aspects (Table 1). For example, Zhao et al. (2024) presented a broad survey of LM interpretability (e.g., generating textual explanations for LMs) but did not focus on MI. Concurrent with our work, Ferrando et al. (2024) and Bereska & Gavves (2024) reviewed techniques developed for understanding the inner workings of transformer-based LMs; however, they present literature from a technique-centric perspective and have a limited scope (e.g., mainly focusing on AI safety and without a focus on the applications of MI). In comparison, we present a unique perspective of a task-centric review and provide a comprehensive scope of MI advancements, aiming to enable beginners of MI to quickly pick up the field and leverage MI in their tasks.
@@
Our paper is organized as follows. We first present the background knowledge needed to understand this survey (Section 2). Following this, we introduce the three fundamental objects of study of MI, which conceptually structure the research explorations in MI Olah et al. (2020) and also discuss the current (ambiguous) uses of the term “MI” in literature (Section 3). Next, we introduce various techniques used in MI research (Section 4). For each technique, we introduce (1) its basic concepts and connections to other sub-fields of interpretability or to work that do not target transformer-based LMs, (2) recent advancements, particularly the limitations of the technique and what has or has not been addressed by further research, and (3) technique-specific evaluations. Then, we introduce Beginner’s Roadmap, which provides a task-centric overview of MI (Section 5). Corresponding to the fundamental objects of study, the roadmap summarizes the actionable workflows along three main lines, i.e., the study of features, the study of circuits, and the study of universality. Subsequently, we present case studies that map the research activities of prior works to the workflow presented in the beginner’s roadmap (Section 6). Next, we present an overview of findings and applications for each study in the roadmap (Section 7). This section includes an in-depth discussion of the findings and novel contributions of various existing works. Finally, we provide careful discussion on the challenges and issues associated with the current development of MI, as well as suggestions for future work (Section 8).
@@ 3 What is Mechanistic Interpretability?
The goal of MI is to reverse engineer the detailed computations performed by a model into human-understandable algorithms, similar to how a programmer might try to reverse engineer complicated binaries into human-readable source code. To achieve this goal, MI takes a bottom-up approach by decomposing the model into smaller components and more elementary computations Zou et al. (2023). By understanding these smaller components and their interactions, MI aims to build a comprehensive understanding of the full model. Below, we first present the three fundamental objects of study in MI, and then discuss its connections and differences with respect to other sub-fields of AI interpretability.
@@ 4 Techniques and Evaluation Methods
In this section, we review major techniques that have been developed to study the fundamental objects described in Section 3.1 for understanding transformer-based LMs. During the introduction of each technique, we will present its basic concepts, technique-specific evaluation, and recent advancements.
@@ 4.1.3 Technical Advancements
Since the introduction of the logit lens, several advancements have been made to enhance its capabilities or to address its limitations, leading to the development of various vocabulary projection methods. Specifically, the advancements can be categorized into the following three categories – (1) Reliability: improving the faithfulness of the technique’s output and enabling more reliable vocabulary projection (2) Decoding position: adapting and applying the technique to decode features from various intermediate representations and model weights; and (3) Decoding expressivity: enhancing the degree to which the technique decodes features beyond immediate next-token predictions.
@@
While the logit lens was initially proposed to decode an LM’s prediction information from activations, subsequent work has applied or adapted the technique to decode information from model weights and gradients. We summarize the various decoding positions as follows. (1) Activations: In addition to the logit lens nostalgebraist (2020), the tuned lens Belrose et al. (2023) and the approach of Din et al. (2023) were both proposed to project the activations of RS, FF and MHA sub-layers into the vocabulary space. In contrast, the attention lens Sakarvadia et al. (2023) was proposed specifically for projecting the output activations of individual attention heads within MHA sub-layers. (2) Model weights: Geva et al. (2022) showed that the logit lens can also be used to interpret the weights of transformer components. Specifically, Geva et al. (2022) used it to understand the role of the FF sublayer in the LM’s prediction by projecting columns of the parameter matrix $W_{v}^{l}$ to the vocabulary space. Dar et al. (2023) further proposed to apply the technique to project the FF sub-layers ($W_{k}^{l}$ and $W_{v}^{l}$) and the query-key ($W_{Q}^{l,h}W_{K}^{l,h}$) and value-output ($W_{V}^{l,h}W_{O}^{l}$) interaction parameter matrices of individual attention heads to the vocabulary space, aiming to study how they transfer and mix information from source tokens to the target token. (3) Gradients: As a representative work, Katz et al. (2024) proposed backward lens, which projects the gradient of FF sub-layers during back-propagation to the vocabulary space, enabling the investigation of how new information is stored in the FF sub-layers of LMs during training.
@@ 4.2.3 Technical Advancements
Intervention-based techniques have undergone several refinements to ensure their correct usage, including various adaptations to address diverse investigative objectives. Specifically, we summarize these technical advancements as follows: (1) Reliability of corruption: ensuring that the corrupt run does not take the model out of distribution (OOD); (2) Intervention for localization: adapting the techniques to identify important components associated with specific LM behavior; (3) Scaling up and automating localization techniques: automating the intervention process to alleviate the need for human effort; (4) Intervention for validating hypotheses: applying intervention to validate a hypothesized interpretation of the LM; and (5) Integrating intervention with vocabulary projection for activation decoding: devising intervention methods for more expressive vocabulary projections of activations.
@@
Figure 4: An example of noising-based intervention on GPT-2 at the node and the edge levels. In the example, a corrupt prompt, with the singular “key” replaced by a plural “keys”, was introduced to discover the circuit in GPT-2 for associating the verb with different forms of the noun. (1) Clean Run: The computational graph of GPT-2 when run on clean input. (2) Corrupt Run: The computational graph of GPT-2 when run on corrupt input. (3) Node Patching: The computational graph of GPT-2 when run on clean input, with the output activation of MHA10 replaced by its activation from the corrupt run (i.e., patching the node of “MHA10”). (4) Edge Patching: The computational graph of GPT-2 when run on clean input, with activations along the edge of (MHA10, FF11) replaced by activations from the corrupt run (i.e., patching the edge of “(MHA10, FF11)”). Colored connections indicate activations calculated from the clean run (Blue), the corrupt run (Red), and the re-calculation after patching (Green).
@@
As we will introduce in our roadmap (Section 5), intervention is commonly used to localize important components (i.e., nodes of circuits) and computational pathways (i.e., edges of circuits) when one tries to discover circuits in an LM. Depending on the goal, the scope of intervention varies. (1) Patching nodes vs. edges: Activation patching is a widely used technique employed to localize key components Meng et al. (2022). This method evaluates the importance of a component $C_{i}$ by patching the activation of $C_{i}$ from the clean run into the corrupt run, one component at a time, to observe its impact on the model’s output, as illustrated in Figure 4. The higher impact implies $C_{i}$ is an important component of the behavior. Path Patching is an extension of activation patching to localize important edges between components Wang et al. (2022a); Goldowsky-Dill et al. (2023). For instance, to assess whether the connection between two components, $C_{1}$ and $C_{2}$, is significant, path patching applies activation patching to the output of $C_{1}$ but, notably, only along paths serving as input to $C_{2}$. In other words, when performing path patching, we allow for the effect of a patched activation only along the computational path we want to investigate, while any other directed nodes will not receive the patched activation despite the connection. Similar to patching nodes, if a change in the LM’s behavior is observed, the connection between the two components is considered important. (2) Patching at various granularity levels: As discussed in Section 3.1.1, it is hypothesized that LMs encode features within a linear subspace of the representation space, potentially represented in a non-standard basis, as indicated by the discovery of polysemantic neurons (see Section 7.1 for further discussion). However, activation patching focuses on patching the entire representation. To address this, Geiger et al. (2024) proposed Distributed Interchange Interventions, also known as Distributed Alignment Search (DAS), which targets interventions within specific subspaces instead of patching full representations.
@@
Performing interventions to localize important components or connections requires humans in the loop, as it involves iteratively patching each component or connection within an LM to assess its significance, which can be time-consuming. To automate this process, Conmy et al. (2023) introduced ACDC (Automatic Circuit DisCovery), which iteratively knocks out edges in the computational graph and removes any edge whose effect on the target metric is less than a specified threshold. However, ACDC is not scalable to large LMs as it requires independent inferences for every iteration. To address the issue, Syed et al. (2023) proposed Edge Attribution Patching (EAP) for locating important edges, where it employs Attribution Patching Nanda et al. (2022) to approximate activation patching for locating important components, which requires only two forward passes and one backward pass for measuring all model components. In addition, Hanna et al. (2024b) recently augmented this approach with Integrated Gradient Sundararajan et al. (2017) to address the concern of zero gradients, leading to the approach of EAP-IG.

### v4

@@ 1 Introduction
Observing these promises and challenges, we aim to provide a comprehensive review of MI in its applications to interpret transformer-based LMs. In particular, our survey offers a task-centric perspective (see Section 2 for a comparison with related surveys), organizing our survey around MI research that investigates fundamental objects of MI study and enabling readers new to the field to quickly identify problems of their interest. Specifically, this is achieved by structuring our taxonomy as a Beginner’s Roadmap to MI (Section 6), where each category and subcategory are centered on overarching research questions and accompanied by actionable workflows. These workflows outline key steps and applicable techniques, guiding practitioners through the entire process from problem formulation to practical implementation. In addition, we discuss the pros and cons of applicable techniques (e.g., vocabulary project methods [nostalgebraist 2020] vs. sparse autonecoders [Bricken et al. 2023]) and present readers with case studies of how prior works fit within the workflow, aiming to aid them in understanding the practices and choosing the best-fit approaches to meet their need. By doing so, our survey serves as a friendly and practical guide for beginning researchers (e.g., new doctoral students) to quickly grasp the field of MI, and for experienced practitioners to more easily organize their study plan.
@@
Our paper is organized as follows. We first discuss how our survey brings unique perspectives and contributions compared to existing surveys on related topics (Section 2), and then present the background knowledge needed to understand this survey (Section 3). Following this, we introduce the three fundamental objects of study of MI, which conceptually structure the research explorations in MI Olah et al. (2020) and also discuss the current (ambiguous) uses of the term “MI” in literature (Section 4). Next, we introduce various techniques used in MI research (Section 5). For each technique, we introduce (1) its basic concepts and connections to other sub-fields of interpretability or to work that do not target transformer-based LMs, (2) recent advancements, particularly the limitations of the technique and what has or has not been addressed by further research, and (3) technique-specific evaluations. Then, we introduce Beginner’s Roadmap , which provides a task-centric overview of MI (Section 6). Corresponding to the fundamental objects of study, the roadmap summarizes the actionable workflows along three main lines, i.e., the study of features, the study of circuits, and the study of universality. Subsequently, we present case studies that map the research activities of prior works to the workflow presented in the beginner’s roadmap (Section 7). Next, we present an overview of findings and applications for each study in the roadmap (Section 8). This section includes an in-depth discussion of the findings and novel contributions of various existing works. Finally, we provide careful discussion on the challenges and issues associated with the current development of MI, as well as suggestions for future work (Section 9).
2 Related Work
Related Survey MI Focused? Format Techniques LM Findings Applications Future Research
Zhao et al. (2024) Technique-centric
Dang et al. (2024) Technique-centric
Ferrando et al. (2024) Technique-centric
Bereska & Gavves (2024) Technique-centric
Shu et al. (2025) (only SAEs) Technique-centric
Sharkey et al. (2025) Task-centric
Ours Task-centric
Table 1: Comparison of our survey with others on LM interpretability. “MI Focused”: whether the survey has a strong focus on MI or whether it includes little or no coverage of MI research; “Format”: whether the survey is elaborated from a technique-centric or task-centric angle (see Section 2 for detail); “Techniques”: whether the survey explains the involved technical approaches in detail (regardless of its MI/non-MI focus); “LM Findings”: whether the survey includes findings and scientific discoveries about LMs’ intrinsic properties by applying the techniques; “Applications”: whether the survey covers how the techniques have been applied to practical tasks; and “Future Research”: whether the survey includes a dedicated section discussing remaining challenges and future work of the field. (*surveys published after our submission.)
There exist surveys on relevant topics, but they differ from ours in a number of key aspects (Table 1). First, our survey focus is exclusively on MI for interpreting LMs, in contrast to some existing surveys Zhao et al. (2024); Dang et al. (2024) that provide a broad overview of LM explainability (e.g., generating textual explanations for LMs). Secondly, among the surveys that focus on MI, one of the key distinctions between the existing surveys and ours is the format or organization of the survey. Specifically, most existing surveys follow an technique-centric format where the central focus is on introducing various interpretability and explainability techniques, and they thus organize the survey around the technique family Zhao et al. (2024); Bereska & Gavves (2024) or their general purpose Ferrando et al. (2024); Dang et al. (2024); Shu et al. (2025). In contrast, our survey adopts a task-centric organization that centers around the practical objectives and workflows involved in MI study. Specifically, the task-centric organization of the survey is enabled by the Beginner’s Roadmap to MI (overview in Section 4.2 and details in Section 6), which (1) organizes the survey around three fundamental objectives of study in MI, i.e., how to perform feature study, circuit study, and the study of universality of features and circuits—which we consider as three core tasks in MI research and practices; (2) elaborates on each task category with its actionable workflows, which distill key steps of the task and suggest appropriate MI techniques for each step; and (3) bridges the workflow concepts and practices through case studies, showing how these workflows have been applied in prior work to solve real MI problems. While technique-centric MI surveys such as Bereska & Gavves (2024) and Ferrando et al. (2024) provide thorough introductions to various MI techniques and their technical developments, their technique-centric format does not include step-by-step workflows and illustrative case studies that can guide readers from research goals to method selection. Our survey addresses this gap by combining technical coverage with actionable guidance. We believe that our task-centric organization provides a more intuitive entry point for newcomers who may feel overwhelmed by the breadth of technical terminology, while also providing experienced practitioners with a structured framework to align their research goals with the most relevant MI techniques.
Beyond its task-centric organization, our survey also differs from other surveys Dang et al. (2024); Shu et al. (2025); Sharkey et al. (2025) in comprehensive coverage of MI techniques. Specifically, we explain the technical details of each technique, outline their limitations and technical advancements, and finally show how they are employed in the step-by-step workflows of the Beginner’s Roadmap to MI. In contrast, Sharkey et al. (2025) and Dang et al. (2024) do not cover the technical introduction of MI techniques, and Shu et al. (2025) is limited to a single MI technique, Sparse Autoencoder (SAE), which we cover in Section 5.3. Similarly, our survey also provides comprehensive discussions on LM findings – how MI techniques have been applied to discover and understand the properties of LMs, such as the functions of their components and their behaviors. This contrasts with surveys such as Zhao et al. (2024) and Dang et al. (2024) that primarily focus on the advancements and comparison of explanation techniques (partially also due to their lack of focus on MI), and Shu et al. (2025) that has only a brief discussion about findings on LM behaviors and is limited to the application of SAEs. In addition, our survey also highlights studies that have applied MI to practical downstream applications such as knowledge editing, LM steering, and AI safety, which were addressed to a lesser extent in Ferrando et al. (2024). Finally, our survey concludes by providing novel discussions on the important challenges associated with the current development of MI and directions for future research, in line with existing surveys.
@@ 4 What is Mechanistic Interpretability?
The goal of MI is to reverse-engineer the detailed computations performed by a model into human-understandable algorithms, similar to how a programmer might try to reverse-engineer complicated binaries into human-readable source code. To achieve this goal, MI takes a bottom-up approach by decomposing the model into smaller components and more elementary computations Zou et al. (2023). By understanding these smaller components and their interactions, MI aims to build a comprehensive understanding of the full model. Below, we first present the three fundamental objects of study in MI, then give an overview of our devised beginner’s roadmap to MI, which summarizes the actionable workflows of current MI studies towards addressing the three fundamental objects, and finally discuss its connections and differences with respect to other sub-fields of AI interpretability.
@@ 4.1.3 Universality
4.2 An Overview of Beginner’s Roadmap to MI
Figure 3: Beginner’s roadmap to MI, designed to help newcomers quickly pick up the field. The MI study is organized into three main categories – feature study, circuit study, and universality study, corresponding to the fundamental three objects of MI study. The roadmap also provides a step-by-step actionable workflow for each category that distills the key steps for completing tasks in that category. More detailed figures and explanations for each category, together with the techniques available at each stage and their respective advantages and disadvantages, are provided in Section 6.
The MI community has explored various approaches to study the three fundamental objects of study. A key motivation for this survey is to provide a friendly guide for researchers and developers interested in MI to quickly pick up the field. To achieve this, we organize our survey in a task-centric format, which is enabled by beginner’s roadmap to MI. We present an overview of the roadmap in Figure 3 and elaborate on its details in Section 6. The roadmap categorizes MI research into three studies, corresponding to the three fundamental objects, i.e., feature study, circuit study, and the study of universality. We provide a general step-by-step workflow for each category. For instance, we propose to further divide feature study into two sub-categories, each with its own distinct workflow: targeted feature study and open-ended feature study. A targeted feature study consists of two steps: Hypothesis generation, in which we propose what feature may be present in the activations, and Hypothesis validation, in which we test whether the feature is indeed represented. In contrast, an open-ended feature study assumes no prior knowledge of what features are encoded and involves two steps: Observe, where we collect information about model activations, and Explain, where we annotate the features encoded based on the information gathered during the Observe step. Similarly, we further categorize circuit study into two sub-categories, each with its own workflow: interpret an LM behavior and interpret an LM component. To interpret an LM behavior, the workflow consists of four steps: Choose LM behavior and dataset; Define computational graph of the selected LM; Localization of the important nodes and edges of the circuit; and Explain circuit components identified during the Localization step. On the other hand, to interpret an LM component, one can begin directly from the Explain circuit components step. Finally, the universality study workflow consists of three steps: Define scope of study, which can be considered along two dimensions—universality of features or universality of circuits; Dimensions of variations, referring to universality across LMs or across tasks; and conduct Feature or circuit study based on the defined scope and variations. Importantly, all three categories also require a final Evaluation step to verify the results of the study. The detailed workflows, together with the techniques available at each stage and their respective advantages and disadvantages, are presented in Section 6. We further provide case studies in Section 7 that illustrate how prior work has applied the proposed roadmap.
@@ 5 Techniques and Evaluation Methods
Techniques Basic Concepts Technique-specific Interpretation Technical Advancements
Vocabulary Projection Methods (Sec 5.1) Decode features from activations by projecting them into the vocabulary space, which is typically achieved by multiplying the activations with the unembedding matrix $W_{U}$, yielding logit distributions over the vocabulary. Infer encoded features by inspecting top/bottom tokens in the projected logit distribution (e.g., a majority of top-k tokens relate to “breakfast” implies a “breakfast” feature). (1) Reliability: improve faithfulness of projections, (2) Decoding positions: enable decoding from activations, weights, and gradients, and (3) Expressivity: decode features beyond next-token predictions.
Intervention-based Methods (Sec 5.2) Investigate LM behavior causally by intervening on intermediate activations during the forward pass and comparing outputs before and after the intervention to infer their role. There are two types: noising-based and denoising-based interventions. Evaluate intervention effects by measuring changes in target token logit, target token probabilities, or the full logit distribution before and after intervention. If an activation is critical, intervening in its value should yield substantial performance changes. (1) Reliability of corruption: improve reliability by ensuring interventions do not take the model out of distribution, (2) Intervention for localization: adapting intervention to discover all behavior-related components, (3) Scaling up and automation: automate the localization technique to alleviate human effort, (4) Intervention for validating hypotheses, and (5) Integrate with vocabulary projection for richer activation decoding.
Sparse Autoencoders (SAE, Sec 5.3) Map activations to a more interpretable but higher-dimensional sparse activations by training SAEs; one can then conduct feature studies in the sparse activations instead of original LM activations. Fidelity of the sparse activations is measured by the reconstruction loss and the interpretability of SAE features is evaluated manually by looking at the activation pattern of the feature or by automated interpretability score. (1) Balancing reconstruction fidelity with sparsity for interpretability, and (2) Discovering features that are functionally useful for downstream LM tasks.
Probing (Sec 5.4.1) Test whether predefined features are present in LM activations by training a shallow classifier (i.e., probe) on them. High probe accuracy on held-out test set indicates feature presence on activations. (Not a focus in this survey; we refer readers to Belinkov (2022) for more details.)
Visualization (Sec 5.4.2) Prepare visualization (e.g., attention maps, neuron activation plots) to investigate model behavior. Human examines visual information to form or refine hypotheses about the roles of components and activations. (Not a focus in this survey.)
Table 2: Summary of MI techniques discussed in Section 5, along with their basic concepts, technique-specific interpretation (i.e., approaches to evaluate the technique outputs), and key technical advancements.
In this section, we review major techniques that have been developed to study the fundamental objects described in Section 4.1 for understanding transformer-based LMs. During the introduction of each technique, we will present its basic concepts, technique-specific interpretation, and recent advancements. A summary is also present in Table 2.
@@ 5.1.3 Technical Advancements
Since the introduction of the logit lens, several advancements have been made to enhance its capabilities or to address its limitations, leading to the development of various vocabulary projection methods. Specifically, the advancements can be categorized into the following three categories – (1) Reliability: improving the faithfulness of the technique’s output and enabling more reliable vocabulary projection (2) Decoding position: adapting and applying the technique to decode features from various intermediate representations and model weights; and (3) Decoding expressivity: enhancing the degree to which the technique decodes features beyond immediate next-token predictions. We highlight representative approaches discussed in the technical advancements in Table 3.
Approach Basic Concepts Technical Advancements
Logit Lens + Translators Belrose et al. (2023); Din et al. (2023); Sakarvadia et al. (2023) Propose to use learned translators for mapping the intermediate activations to the final-layer representation space before applying the logit lens. Improve the reliability of logit lens projection.
Attention Lens Sakarvadia et al. (2023) Interpret attention head outputs with the logit lens and learned head-specific translators. Adapt the logit lens to new decoding positions (i.e., activations of attention heads).
Geva et al. (2022) Showcase that the logit lens can also be used to decode the knowledge stored in the FF sub-layer by projecting the columns of the parameter matrix $W_{v}^{l}$ in vocabulary space. Adapt the logit lens to new decoding positions (i.e., model weights).
Backward Lens Katz et al. (2024) Investigate the knowledge updates in FF neurons during model training by projecting their gradients into the vocabulary space. Adapt the logit lens to new decoding positions (i.e., model gradients).
Attribute Lens Hernandez et al. (2023) Decode objects encoded for a relation from subject activations (e.g., “Space Needle — is located in → Seattle”) by learning a relation-specific linear map and using $W_{U}$ to project in vocabulary space. Improve expressivity of technique beyond decoding next-token prediction.
Future Lens Pal et al. (2023) Analyze whether an activation at timestep $i$ encodes future output tokens at $i+n,n\geq 2$, with two types of approaches: learning a linear mapping from $h_{i}^{l}$ to $h_{i+n}^{L}$ for projection, and patching $h_{i}^{l}$ to a separate LM under a pre-defined or learnable investigation prompt. Improve expressivity of technique beyond decoding next-token prediction.
Table 3: Representative approaches discussed in the technical advancements of vocabulary projection methods.
@@
While the logit lens was initially proposed to decode an LM’s prediction information from activations, subsequent work has applied or adapted the technique to decode information from model weights and gradients. We summarize the various decoding positions as follows. (1) Activations: In addition to the logit lens nostalgebraist (2020), the tuned lens Belrose et al. (2023) and the approach of Din et al. (2023) were both proposed to project the activations of RS, FF, and MHA sub-layers into the vocabulary space. In contrast, the attention lens Sakarvadia et al. (2023) was proposed specifically for projecting the output activations of individual attention heads within MHA sub-layers. (2) Model weights: Geva et al. (2022) showed that the logit lens can also be used to interpret the weights of transformer components. Specifically, Geva et al. (2022) used it to understand the role of the FF sublayer in the LM’s prediction by projecting columns of the parameter matrix $W_{v}^{l}$ to the vocabulary space. Dar et al. (2023) further proposed to apply the technique to project the FF sub-layers ($W_{k}^{l}$ and $W_{v}^{l}$) and the query-key ($W_{Q}^{l,h}W_{K}^{l,h}$) and value-output ($W_{V}^{l,h}W_{O}^{l}$) interaction parameter matrices of individual attention heads to the vocabulary space, aiming to study how they transfer and mix information from source tokens to the target token. (3) Gradients: As a representative work, Katz et al. (2024) proposed backward lens, which projects the gradient of FF sub-layers during back-propagation to the vocabulary space, enabling the investigation of how new information is stored in the FF sub-layers of LMs during training.
@@ 5.2.3 Technical Advancements
Intervention-based techniques have undergone several refinements to ensure their correct usage, including various adaptations to address diverse investigative objectives. Specifically, we summarize these technical advancements as follows: (1) Reliability of corruption: ensuring that the corrupt run does not take the model out of distribution (OOD); (2) Intervention for localization: adapting the techniques to identify important components associated with specific LM behavior; (3) Scaling up and automating localization techniques: automating the intervention process to alleviate the need for human effort; (4) Intervention for validating hypotheses: applying intervention to validate a hypothesized interpretation of the LM; and (5) Integrating intervention with vocabulary projection for activation decoding: devising intervention methods for more expressive vocabulary projections of activations. We present representative approaches discussed in the technical advancements in Table 5.
Approach Basic Concepts Technical Advancements
Mean-ablation Rai & Yao (2024) Replace activation with its average across counterfactual inputs, preserving baseline task information while removing or corrupting input-specific contributions. Improve reliability of corruption by mitigating the OOD concern due to intervention.
Resampling-ablation Rai & Yao (2024) Replace activation with its activation of a randomly sampled counterfactual input. Improve reliability of corruption by mitigating the OOD concern due to intervention.
Activation Patching Meng et al. (2022) Investigate the importance of a node in the circuit by performing noising or denoising intervention on the node activation. Intervention for localizing an important node in a circuit.
Path Patching Wang et al. (2022a) Investigate the importance of edges between two components ($C_{1}$ and $C_{2}$) by applying activation patching only along the computational path from $C_{1}$ to $C_{2}$. Intervention for localizing important edges in a circuit.
Distributed Interchange Interventions (DII) Geiger et al. (2024) Target interventions to specific subspaces of a representation rather than replacing the entire activation. Intervention at a more granular level than activation patching.
Distributed Alignment Search (DAS) Geiger et al. (2024) A supervised method for discovering lower-dimensional subspaces associated with specific causal variables. Intervention at a more granular level than activation patching.
ACDC Conmy et al. (2023) Automate circuit discovery by iteratively removing edges in the computational graph and discarding those with minimal effect on a target metric. Automate intervention for localization by reducing manual effort.
EAP Syed et al. (2023) & EAP-IG Hanna et al. (2024b) Approximate edge patching with gradient-based attribution to efficiently locate important edges in a circuit, trading some faithfulness for much greater scalability. Automate and scale up intervention for localization by reducing the computational requirement.
Causal Scrubbing Chan et al. (2022a) Rigorously test a hypothesis about how a model works by systematically intervening in its internal activations and then observing if the outputs (before and after intervention) match what the hypothesis would predict. Intervention for validating hypotheses.
Fixed/Learned Prompt Causal Intervention Pal et al. (2023) Decode features from intermediate activations by patching them into a separate LM run on a fixed or learned prompt, then interpreting the resulting generation as the features encoded in the original activation. The initial goal of the approach is to decode future tokens from the end of the input. Integrate intervention with vocabulary projection for activation decoding.
Patchscope Ghandeharioun et al. (2024) Decode features from intermediate activations by patching them from a source LM and prompt into a target LM run on an inspection (target) prompt, then interpreting the resulting generation as the features encoded in the original activation. Integrate intervention with vocabulary projection for activation decoding.
Table 5: Representative approaches in the technical advancements of intervention techniques.
@@
Figure 5: An example of noising-based intervention on GPT-2 at the node and the edge levels. In the example, a corrupt prompt, with the singular “key” replacing the plural “keys” in the clean prompt, was introduced to discover the circuit in GPT-2 for associating the verb with different forms of the noun. (1) Clean Run: The computational graph of GPT-2 when run on clean input. (2) Corrupt Run: The computational graph of GPT-2 when run on corrupt input. (3) Node Patching: The computational graph of GPT-2 when run on clean input, with the output activation of MHA10 replaced by its activation from the corrupt run (i.e., patching the node of “MHA10”). (4) Edge Patching: The computational graph of GPT-2 when run on clean input, with activations along the edge of (MHA10, FF11) replaced by activations from the corrupt run (i.e., patching the edge of “(MHA10, FF11)”). Colored connections indicate activations calculated from the clean run (Blue), the corrupt run (Red), and the re-calculation after patching (Green), respectively.
@@
As we will introduce in our roadmap (Section 6), intervention is commonly used to localize important components (i.e., nodes of circuits) and computational pathways (i.e., edges of circuits) when one tries to discover circuits in an LM. Depending on the goal, the scope of intervention varies. (1) Patching nodes vs. edges: Activation patching is a widely used technique employed to localize key components Meng et al. (2022). This method evaluates the importance of a component $C_{i}$ by patching the activation of $C_{i}$ from the clean run into the corrupt run, one component at a time, to observe its impact on the model’s output, as illustrated in Figure 5. The higher impact implies $C_{i}$ is an important component of the behavior. Path Patching is an extension of activation patching to localize important edges between components Wang et al. (2022a); Goldowsky-Dill et al. (2023). For instance, to assess whether the connection between two components, $C_{1}$ and $C_{2}$, is significant, path patching applies activation patching to the output of $C_{1}$ but, notably, only along paths serving as input to $C_{2}$. In other words, when performing path patching, we allow for the effect of a patched activation only along the computational path we want to investigate, while any other directed nodes will not receive the patched activation despite the connection. Similar to patching nodes, if a change in the LM’s behavior is observed, the connection between the two components is considered important. (2) Patching at various granularity levels: As discussed in Section 4.1.1, it is hypothesized that LMs encode features within a linear subspace of the representation space, potentially represented in a non-standard basis, as indicated by the discovery of polysemantic neurons (see Section 8.1 for further discussion). However, activation patching focuses on replacing the entire hidden representation and is typically analyzed under a localist assumption that each high-level causal variable corresponds to a disjoint set of neurons (Geiger et al., 2024; 2025). To overcome this limitation, Distributed Interchange Interventions (DII) proposes to perform intervention in rotated subspaces of the representation rather than in the raw neuron basis (Geiger et al., 2024). Concretely, DII performs the following operations: (i) it first applies a change-of-basis rotation transformation to rotate the model’s representation into a new basis; (ii) within this rotated basis, specific subspaces corresponding to high-level causal variables are identified, and interventions are performed on those subspaces; (iii) after the intervention, the representation is transformed back into the original neural basis. By operating at the subspace level, DII reveals interpretable distributed structure and enables more fine-grained interventions of model activations. In addition to DII, Geiger et al. (2024) propose another technique Distributed Alignment Search (DAS), a supervised method for automatically discovering the rotation matrix and the $k$-dimensional subspaces within the rotated representation that best align with the high-level causal variables. DAS involves learning an orthogonal transformation matrix that maps neural representations onto $k$-dimensional subspaces aligned with the causal variable. The transformation is trained on clean-counterfactual pairs via interchange interventions, with the objective of maximizing interchange intervention accuracy under the $k$-dimensional constraint.
@@
Performing interventions to localize important components or connections requires humans in the loop, as it involves iteratively patching each component or connection within an LM to assess its significance, which can be time-consuming. To automate this process, Conmy et al. (2023) introduced ACDC (Automatic Circuit DisCovery), which iteratively knocks out edges in the computational graph and removes any edge whose effect on the target metric is less than a specified threshold. However, ACDC is not scalable to large LMs as it requires independent inferences for every iteration. To address the issue, Syed et al. (2023) proposed Edge Attribution Patching (EAP) for locating important edges, where it employs Attribution Patching Nanda et al. (2022) to approximate activation patching for locating important edges, which requires only two forward passes and one backward pass for measuring all model components. In addition, Hanna et al. (2024b) recently augmented this approach with Integrated Gradient Sundararajan et al. (2017) to address the concern of zero gradients, leading to the approach of EAP-IG.